OpenStudy (anonymous):

solve for x: (x^x^2) * (2x * (ln(x)) + x) = 0 Please show steps, thanks! solve for x: (x^x^2) * (2x * (ln(x)) + x) = 0 Please show steps, thanks! @Mathematics

OpenStudy (turingtest):

\[x^{x^2} (2x \ln(x) + x) = 0\] right?

OpenStudy (anonymous):

yeah that's right

OpenStudy (anonymous):

it's the derivative of \[x^{x^{2}}\]

OpenStudy (anonymous):

trying to find the interval which the derivative is >0 and >0. Having some trouble solving this though.

OpenStudy (turingtest):

I see Wolfram has the answer\[ x={1 \over \sqrt e}\]but it cannot show steps, I will continue to work on it and get back to you if I have any progress.

OpenStudy (anonymous):

ok, thanks

OpenStudy (turingtest):

Trying to find an interval which is what? Has derivative >0?

OpenStudy (anonymous):

trying to find the intervals where the derivative is greater than 0 and less than 0

OpenStudy (anonymous):

ok I've got it, you just multiply both sides by the (x^(x^2)) and just solve 2x * (ln(x)) + x = 0